# Kolmogorov-Arnold Attention: Is Learnable Attention Better For Vision Transformers?

### v1

@@ 3 How Can We Design Learnable Attention?
In this Section, we design learnable multi-head self-attention (MHSA) module for ViTs that operate inside its encoder blocks.
Last year witnessed a surge in embedding KANs in different DNN architectures (Genet & Inzirillo, 2024; Ferdaus et al., 2024; Li et al., 2024). To our knowledge, the work closely related to ours is KAT, where (Yang & Wang, 2025) replace the MLP layers in the ViTs with KAN layers; also, see (Chen et al., 2024). We later show that deploying learnable attention is a complicated and much-involved process.
@@
Let ${\cal A}^{i,j}\in\mathbb{R}^{N\times N}$ be the attention matrix for $i^{\rm th}$ head in the $j^{\rm th}$ encoder block. For $k\in[N],$ the softmax activation function, $\sigma:\mathbb{R}^{N}\to(0,1)^{N}$, operates on the $k^{\rm th}$ row vector, ${\cal A}^{i,j}_{k,:}\in\mathbb{R}^{1\times N}$ of ${\cal A}^{i,j}$ to produce component-wise output $\sigma({\cal A}^{i,j}_{k,l})=\tfrac{e^{{\cal A}^{i,j}_{k,l}}}{\sum\limits_{n=1}^{N}e^{{\cal A}^{i,j}_{k,n}}}.$ Instead of using the softmax function, we can use a learnable activation function, ${\tilde{\sigma}}$ on the row vectors of each attention head ${\cal A}^{i,j}$. With any choice of the basis functions (e.g., B-Spline, Fourier, Fractals, Wavelets, etc.), the activated attention row vector, ${\tilde{\sigma}(\mathcal{A}^{i,j}_{k,:})}$ for $k\in[N]$ can be written as
@@
$\textstyle{\tilde{\sigma}\left[({\cal A}^{i,j}_{k,:})\right]}$ $=\textstyle{\begin{pmatrix}\phi_{11}(\cdot)&\phi_{12}(\cdot)&...&\phi_{1N}(\cdot)\\ \phi_{21}(\cdot)&\phi_{22}(\cdot)&...&\phi_{2N}(\cdot)\\ \vdots&\vdots&\ddots&\vdots\\ \phi_{N1}(\cdot)&\phi_{N2}(\cdot)&...&\phi_{NN}(\cdot)\end{pmatrix}}\scriptscriptstyle{\begin{pmatrix}{\cal A}^{i,j}_{k,1}\\ {\cal A}^{i,j}_{k,2}\\ \vdots\\ {\cal A}^{i,j}_{k,N}\end{pmatrix}}$
$=\textstyle{\left(\Phi^{i,j}\left[({\cal A}^{i,j}_{k,:})^{\top}\right]\right)^{\top}},$ (1)
@@
where $\Phi^{i,j}=[\phi^{i,j}_{pq}]\in\mathbb{R}^{N\times N}$ and each matrix entry, $\phi_{pq}$, is referred to as a learnable unit, represented using a set of basis functions $\{\psi^{i,j}_{m}\}_{m=1}.$ The coefficients associated with the basis functions are called the learnable parameters. Note that, in our convention, ${{\cal A}^{i,j}_{k,:}\in\mathbb{R}^{1\times N}}$ is a row vector. We apply the transpose operation to each row vector, to adopt the convention used in KAN layers. As a result, ${\Phi^{i,j}\left[({\cal A}^{i,j}_{k,:})^{\top}\right]\in\mathbb{R}^{N\times 1},}$ and we transpose it to obtain learnable attention row vectors ${\tilde{\sigma}\left[({\cal A}^{i,j}_{k,:})\right]}$; see Figures 2(b)-(c).
@@
Projection on to the probability simplex. The softmax function acts on each row of the attention matrix to create a probability distribution; the learnable attention does not ensure that. To ensure that each row vector of the learned attention matrix lies on a probability simplex, we project them onto the $\ell_{1}$-unit ball. That is, for each attention matrix, ${\cal A}^{i,j}\in\mathbb{R}^{N\times N}$, and for each $k\in[N],$ we want to have ${\|\tilde{\sigma}\left[({\cal A}^{i,j}_{k,:})\right]\|_{1}=1}$ and $\tilde{\sigma}\left[({\cal A}^{i,j}_{k,l})\right]\geq 0$ for $l\in[N]$. In the following, we cast it as a sparse approximation problem, whose variants have been well-studied in the past decade and arise frequently in signal processing (Bryan & Leise, 2013; Donoho, 2006; Dutta, 2016) and matrix approximation (Boas et al., 2017; Stewart, 1993).
A sparse approximation problem. Let $y\in\mathbb{R}^{m}$ be a given vector. If we want to approximate $y$ with a vector, $x\in\mathbb{R}^{m}$, with positive components and $\|x\|_{1}=1$, we can write the constrained optimization problem as:
${}x^{\star}=\arg\min_{x\in\mathbb{R}^{m}}\frac{1}{2}\|x-y\|^{2}$
$\text{subject to\;\;}x_{i}\geq 0\;\text{for}\;i\in[m],\text{and}\;\|x\|_{1}=1.$ (2)
For completeness, we describe the solution to problem (3) in §A. Algorithm 1 provides a pseudocode for $\ell_{1}$ projection, and by setting $y=\tilde{\sigma}\left[({\cal A}^{i,j}_{k,:})\right]$ and $m=N$ for each $k\in[N]$ in Algorithm 1, we obtain the projected attention vector.
@@ 4 Our Architecture
Section 3 outlines a generic learnable attention module. However, we experience computation bottleneck in implementing it. As in (3), for learnable activation function, $\tilde{\sigma}:\mathbb{R}^{N}\to\mathbb{R}^{N}$, an operator matrix, $\Phi^{i,j}\in\mathbb{R}^{N\times N}$ acts on each row ${\cal A}^{i,j}_{k,:}$ of ${\cal A}^{i,j}$, and produces the output $\tilde{\sigma}({\cal A}^{i,j}_{k,:}).$ These operations are extremely compute-heavy and have large memory footprints. E.g., ViT-tiny (Dosovitskiy et al., 2020; Wu et al., 2022; Steiner et al., 2022) has 5.53M parameters and requires nearly 0.9 GB of GPU memory to train. Implementing the learnable attention (3) in ViT-tiny training with B-splines of order 5 and grid size 10 and Fourier basis with grid size 10 increases the parameter count to 30.68M and 39.06M, respectively. We could not train the version with B-splines for its humongous memory requirements, and B-spline computations are non-parallelizable. The one with Fourier basis is parallelizable but it takes approximately 60 GB of GPU memory when computing with a batch of one image. The computational bottleneck is agnostic of the basis functions.
(a) Attention matrix ${\mathcal{A}^{i,j}}$ before softmax activation.
Figure 3: Spectral analysis of the attention matrices before and after softmax shows that they are low-rank. For this experiment, we use all 3 heads in the last encoder block of ViT tiny on 5 randomly sampled images from the CIFAR-10 validation set. We plot all 15 singular vectors (each of 197 dimensions) where the singular values are arranged in non-increasing order.
How can we remedy this? Deep neural networks exhibit low-rank structure (Oja, 1982; Jain et al., 2016). Recently, (Kwon et al., 2024) showed across various networks, the training largely occurs within a low-dimensional subspace. Therefore, we postulate that the attention matrices also have an underlying low-rank structure.
To validate this, we perform spectral analysis of the attention matrices before and after the softmax activation; see Figures 3-3. The scree test (Cattell, 1966) shows that for vanilla attention without softmax, there are 8 significant singular values; after softmax activation this count increases to 16. The above observation verifies that with or without the nonlinear activation, the attention matrix $\mathcal{A}^{i,j}$ possesses an underlying low-rank structure. This motivates us to use a lower-dimensional operator, ${\widehat{\Phi}}$ for the learnable attention calculation with $r=12$. In §B.3, we perform an ablation to find the best $r.$
Instead of using an ${N\times N}$ operator, ${\Phi^{i,j}}$, we use a reduced sized operator ${\widehat{\Phi}^{i,j}\in\mathbb{R}^{r\times N}}$ such that ${r\ll N}$, and the new learned activation is $r$-dimensional for $k\in[N]$. That is, ${\widehat{\Phi}^{i,j}}$ down-projects each attention row vector of ${\cal A}^{i,j}$ to a lower dimensional subspace. This process significantly reduces the computational overhead. Next, we post-multiply using another learnable weight matrix, $W^{i,j}\in\mathbb{R}^{N\times r}$ to project them back to their original dimension. For each $k\in[N]$, this operation results in computing ${\widehat{\sigma}({\cal A}^{i,j}_{k,:})=\left[W^{i,j}\widehat{\Phi}^{i,j}\left[({\cal A}^{i,j}_{k,:})^{\top}\right]\right]^{\top}}$; see Figure 2(d).
@@
Figure 4: Different configurations to update $\widehat{\Phi}$:(a) Blockwise configuration, where $\Phi^{i,1}\neq\Phi^{i,2}\neq\cdots\neq\Phi^{i,L}$ for all $i=1,2,...,h$ (b) universal configuration, where $\Phi^{i,1}=\Phi^{i,2}=\cdots=\Phi^{i,L}=\Phi^{i}$ for all $i=1,2,...,h.$
@@
Fourier Kolmogorov-Arnold Attention. Although the default basis for KANs is B-Splines in (Liu et al., 2025), the number of MHSA parameters is much more than that of an MLP. Specifically, for $L$ encoder blocks, each with $h$ attention heads, the parameter complexity is $O(N^{2}Lh(G+k))$, for $\Phi\in\mathbb{R}^{N\times N}$ and $O(NrLh(G+k))$, for $\widehat{\Phi}\in\mathbb{R}^{r\times N}$, if the model uses B-Splines of degree $k$.
@@
Moreover, as (Yang & Wang, 2025) mentioned, B-splines are localized functions, not standard CUDA functions. Although efficient CUDA implementations for cubic B-Splines exist (Ruijters & Thévenaz, 2012; Ruijters et al., 2008), their overall implementation results in slower, sparse, non-scalable, and complicated GPU execution; also, see (Xu et al., 2024; Pal & Das, 2024). In §B.7, our extensive experiments with B-spline basis in KAN for image classification tasks verify their poor generalizability on medium-scale datasets (e.g., CIFAR-10 and CIFAR-100).
@@
(Carleson, 1967) Let $f$ be an ${L^{p}}$ periodic function for ${p\in(1,\infty]}$, with Fourier coefficients ${\hat{f}(n)}$. Then ${\lim_{N\to\infty}\sum_{|n|\leq N}{\hat{f}}(n)e^{inx}=f(x)},$ for almost every $x$.
In particular, if a function is continuously differentiable, its Fourier series converges to it everywhere. In the past, (Xu et al., 2024; Mehrabian et al., 2024) used Fourier basis in KAN, (Dong et al., 2024) proposed Fourier analysis network (FAN) to capture periodic phenomena. Motivated by these, we use the Fourier basis to approximate the effect of the smooth softmax function. For our architecture, we employ the Fourier basis, $\{\sin{(\cdot)},\cos{(\cdot)}\}$ with gridsize $G$ to design learnable attention. Algorithm 1 can be used optionally. For the operator ${\widehat{\Phi}^{i,j}=[\widehat{\phi}^{i,j}_{pq}]\in\mathbb{R}^{r\times N}}$, we have ${\widehat{\Phi}^{i,j}:\mathbb{R}^{N}\to\mathbb{R}^{r}}$. Hence, each row ${{\cal A}^{i,j}_{k,:}}$, with ${k\in[N]}$, transformed into ${\widehat{\Phi}^{i,j}_{p}\left[({\cal A}^{i,j}_{k,:})^{\top}\right]=\sum_{q=1}^{N}\widehat{\phi}^{i,j}_{pq}({\cal A}^{i,j}_{k,q})},$ via the Fourier bases such that
${\widehat{\phi}^{i,j}_{pq}({\cal A}^{i,j}_{k,q})=\sum_{m=1}^{G}a_{pqm}\cos{(m{\cal A}^{i,j}_{k,q})}+b_{pqm}\sin{(m{\cal A}^{i,j}_{k,q})},}$
where $m$ refers to the ${m^{\rm th}}$ grid point on a uniform grid of size $G$. The weights, [{${[\{[a_{pqm}],[b_{pqm}]\}_{m=1}^{G},W^{i,j}]}$ are updated in the backpropagation. If we use a Fourier basis in ${\widehat{\Phi}\in\mathbb{R}^{r\times N}}$, the parameter complexity of MHSA becomes ${O(2NrhGL)}.$ Regardless of the basis used, the complexity of linear projection is $O(N^{2}rhL).$ We note that for any Transformer, $L$ and $h$ are fixed. So, one may omit them from the complexity results.
Figure 5: Training loss and test accuracy of vanilla ViTs and their Fourier KArAt versions on CIFAR-10, CIFAR-100 and ImageNet-1K.
Blockwise and Universal operator configuration. We consider two configurations for updating the operator $\widehat{\Phi}^{i,j}$. (a) Blockwise: In this configuration, each encoder block involves learning the attention through $h$ distinct operators $\widehat{\Phi}$ for each of the $h$ heads, totaling $hL$ operators; see Figure 4 (a). Like the MHSA architecture in ViTs, the blockwise configuration is designed to learn as many different data representations as possible. (b) Universal: The motivation behind this configuration comes from the KAT (Yang & Wang, 2025). In KAT, the MLP head is replaced with different variations of a KAN head—KAN and Group-KAN. In Group-KAN, the KAN layer shares rational base functions and their coefficients among the edges. Inspired by this, in our update configuration, all $L$ encoder blocks share the same $h$ operators; $\widehat{\Phi}^{i,j}=\widehat{\Phi}^{i}$ for $j=1,2,\ldots,L$; see Figure 4 (b). Rather than learning attention through $hL$ operators, this configuration only uses $h$ operators. Here, we share all learnable units and their parameters in each head across $L$ blocks. We postulate that blockwise mode with more operators captures more nuances from the data. In contrast, the universal mode is suitable for learning simpler decision boundaries from the data. We also note that these two modes can be used with (3) as shown in Figure  2. Finally, Algorithm 2 in Appendix gives pseudocode of applying Fourier-KArAt in ViT encoder block.
Table 1: Performance of the best-performing Fourier KArAt models compared to the conventional vanilla baselines. The best and the second best Top-1 accuracies are given in red and blue, respectively. The $\downarrow$ and $\uparrow$ arrows indicate the relative loss and gain, respectively, compared to the base models.
@@
5 Empirical Analysis
Baselines and Datasets. For our benchmarking, we chose three popular vision Transformers—ViT-tiny, ViT-small (Wu et al., 2022; Steiner et al., 2022), and ViT-base (Dosovitskiy et al., 2020). We incorporate Fourier-KArAt in them by replacing their softmax function with learnable activation. We perform our benchmarking on CIFAR-10, CIFAR-100 (Krizhevsky et al., 2009) and ImageNet-1K (Deng et al., 2009) datasets.

### v2

@@ 3 How Can We Design Learnable Attention?
Last year, we witnessed a surge in embedding KANs in different DNN architectures [29, 28, 44]. To our knowledge, the work closely related to ours is KAT, where [80] replaces the MLP layers in the ViTs with KAN layers; also, see [17]. We note that, for attentive graph neural networks (GNNs), [27] unifies the scoring functions of attentive GNNs and names it as Kolmogorov-Arnold Attention (KAA). However, this attention is orthogonal to our MHSA design; deploying learnable MHSA is complicated. In this Section, we design learnable multi-head self-attention (MHSA) module for ViTs that operate inside its encoder blocks.
@@
Let ${\cal A}^{i,j}\in\mathbb{R}^{N\times N}$ be the attention matrix for $i^{\rm th}$ head in the $j^{\rm th}$ encoder block. Instead of using the softmax function row-wise, we can use a learnable activation function, ${\tilde{\sigma}}$ on the row vectors of each attention head ${\cal A}^{i,j}$. With any choice of the basis functions (e.g., B-Spline, Fourier, Wavelets, etc.), the activated attention row vector, ${\tilde{\sigma}(\mathcal{A}^{i,j}_{k,:})}$ for $k\in[N]$ can be written as
@@
$\textstyle{\tilde{\sigma}\left[({\cal A}^{i,j}_{k,:})\right]}$ $=\textstyle{\begin{pmatrix}\phi_{11}(\cdot)&\phi_{12}(\cdot)&...&\phi_{1N}(\cdot)\\ \phi_{21}(\cdot)&\phi_{22}(\cdot)&...&\phi_{2N}(\cdot)\\ \vdots&\vdots&\ddots&\vdots\\ \phi_{N1}(\cdot)&\phi_{N2}(\cdot)&...&\phi_{NN}(\cdot)\end{pmatrix}}\scriptscriptstyle{\begin{pmatrix}{\cal A}^{i,j}_{k,1}\\ {\cal A}^{i,j}_{k,2}\\ \vdots\\ {\cal A}^{i,j}_{k,N}\end{pmatrix}}=\textstyle{\left(\Phi^{i,j}\left[({\cal A}^{i,j}_{k,:})^{\top}\right]\right)^{\top}},$ (2)
@@
where $\Phi^{i,j}=[\phi^{i,j}_{pq}]\in\mathbb{R}^{N\times N}$ and each matrix entry, $\phi_{pq}$, is referred to as a learnable unit, represented using a set of basis functions $\{\psi^{i,j}_{m}\}_{m=1}.$ The coefficients associated with the basis functions are the learnable parameters. In our convention, ${{\cal A}^{i,j}_{k,:}\in\mathbb{R}^{1\times N}}$ is a row vector. We apply the transpose operation to each row vector to adopt the convention used in KAN layers. Hence ${\Phi^{i,j}\left[({\cal A}^{i,j}_{k,:})^{\top}\right]\in\mathbb{R}^{N\times 1},}$ and we transpose it to obtain learnable attention row vectors ${\tilde{\sigma}\left[({\cal A}^{i,j}_{k,:})\right]}$; see Figures 1(b)(ii)-(iii).
@@
Projection onto the probability simplex. The softmax function acts on each row of ${\cal A}$ to create a probability distribution; the learnable attention does not guarantee that. To ensure that each row vector of $\Phi(\mathcal{A})$ lies on a probability simplex, we project them onto the $\ell_{1}$-unit ball. That is, for each attention matrix, ${\cal A}^{i,j}\in\mathbb{R}^{N\times N}$, and for each $k\in[N],$ we want to have ${\|\tilde{\sigma}\left[({\cal A}^{i,j}_{k,:})\right]\|_{1}=1}$ and $\tilde{\sigma}\left[({\cal A}^{i,j}_{k,l})\right]\geq 0$ for $l\in[N]$. We cast a sparse approximation problem, whose variants have been well-studied in the past decade and arise frequently in signal processing [12, 23, 26] and matrix approximation [8, 69]; see §4. Algorithm 1 provides an $\ell_{1}$-projection pseudocode; for each $k\in[N]$, setting ${y=\tilde{\sigma}\left[({\cal A}^{i,j}_{k,:})\right]}$ and $m=N$ in Algorithm 1, we obtain the projected attention vector.
@@ 4 Our Architecture
Different from learning activations for MLP, as in KAN and KAT, implementing the generic learnable attention incurs impractical computational and memory costs on present computing hardware and DL toolkits. For learnable activation, $\tilde{\sigma}:\mathbb{R}^{N}\to\mathbb{R}^{N}$, a full-rank operator, ${\Phi^{i,j}\in\mathbb{R}^{N\times N}}$ acts on each row ${\cal A}^{i,j}_{k,:}$ of ${\cal A}^{i,j}$, and produces $\tilde{\sigma}({\cal A}^{i,j}_{k,:}).$ These operations are extremely compute-heavy and have large memory footprints. E.g., ViT-Tiny [24, 76, 68] has 5.53M parameters and requires nearly 0.9 GB of GPU memory to train. Implementing the learnable attention (2) in ViT-Tiny training with B-splines of order 5 and grid size 10 and Fourier basis with grid size 10 increases the parameter count to 30.68M and 39.06M, respectively. We could not train the version with B-splines for its humongous memory requirements, and B-spline computations are non-parallelizable. The one with the Fourier basis is parallelizable, but it takes approximately 60 GB of GPU memory when computing with a batch of one image. This computational bottleneck is agnostic of the basis functions.
@@
Figure 2: Different configurations to update $\widehat{\Phi}$:(a) Blockwise configuration, where $\Phi^{i,1}\neq\Phi^{i,2}\neq\cdots\neq\Phi^{i,L}$ for all $i=1,2,...,h$; (b) universal configuration, where $\Phi^{i,1}=\Phi^{i,2}=\cdots=\Phi^{i,L}=\Phi^{i}$ for all $i=1,2,...,h.$
How can we remedy this? Deep neural networks exhibit low-rank structure [60, 36]. Recently, [41] showed that across various networks, the training largely occurs within a low-dimensional subspace. Therefore, we postulate that the attention matrices also have an underlying low-rank structure.
To validate this, we perform spectral analysis of the attention matrices before and after the softmax activation; see Figures 6-6. The scree test [16] shows that for traditional attention matrices without softmax, there are 8 significant singular values; after softmax activation, this count increases to 16. The above observation verifies that with or without the nonlinear activation, the attention matrix $\mathcal{A}^{i,j}$ possesses an underlying low-rank structure. This motivates us to use a lower-dimensional operator, ${\widehat{\Phi}}$ for the learnable attention calculation with $r=12$. In §B.3.2-B.3.3, we perform an ablation to find the best $r.$
Instead of ${\Phi^{i,j}\in\mathbb{R}^{N\times N}}$, we use an operator, ${\widehat{\Phi}^{i,j}\in\mathbb{R}^{r\times N}}$ such that ${r\ll N}$, and the new learned activation is $r$-dimensional for $k\in[N]$. This process significantly reduces the computational overhead. Next, we post-multiply another learnable weight matrix, $W^{i,j}\in\mathbb{R}^{N\times r}$ to project them back to their original dimension. For each ${k\in[N]}$, this operation results in computing ${\widehat{\sigma}({\cal A}^{i,j}_{k,:})=\left[W^{i,j}\widehat{\Phi}^{i,j}\left[({\cal A}^{i,j}_{k,:})^{\top}\right]\right]^{\top}}$ (Figure 1(b)(iv)); see other variants in §B.4. Our approximation is not unique or the best; it is one of the feasible solutions given the present computing interface.
@@
Fourier Kolmogorov-Arnold Attention. Although the default basis for KANs is B-Splines in [51], the number of MHSA parameters is much more than that of an MLP. Specifically, for $L$ encoder blocks, each with $h$ attention heads, the parameter complexity is $O(N^{2}Lh(G+k))$, for $\Phi\in\mathbb{R}^{N\times N}$ and $O(NrLh(G+k))$, for $\widehat{\Phi}\in\mathbb{R}^{r\times N}$, if the model uses B-Splines of degree $k$.
@@
Moreover, as [80] mentioned, B-splines are localized functions, not standard CUDA functions. Although efficient CUDA implementations for cubic B-Splines exist [66, 65], their overall implementation results in slower, sparse, non-scalable, and complicated GPU execution; also, see [78, 61]. In §B.3.1, our extensive experiments with B-spline basis in KAN for image classification tasks verify their poor generalizability on medium-scale datasets (e.g., CIFAR-10 and CIFAR-100).
@@
[14] Let $f$ be an ${L^{p}}$ periodic function for ${p\in(1,\infty]}$, with Fourier coefficients ${\hat{f}(n)}$. Then ${\lim_{N\to\infty}\sum_{|n|\leq N}{\hat{f}}(n)e^{inx}=f(x)},$ for almost every $x$.
In particular, if a function is continuously differentiable, its Fourier series converges to it everywhere. In the past, [78, 53] used the Fourier basis in KAN, [22] proposed a Fourier analysis network (FAN) to capture periodic phenomena. Motivated by these, we use the Fourier basis to approximate the effect of the smooth softmax function. We employ the Fourier basis, $\{\sin{(\cdot)},\cos{(\cdot)}\}$ with gridsize $G$ to design learnable attention. Algorithm 1 can be used optionally. For the operator ${\widehat{\Phi}^{i,j}=[\widehat{\phi}^{i,j}_{pq}]\in\mathbb{R}^{r\times N}}$, we have ${\widehat{\Phi}^{i,j}:\mathbb{R}^{N}\to\mathbb{R}^{r}}$. Hence, each row ${{\cal A}^{i,j}_{k,:}}$, for ${k\in[N]}$, transformed into ${\widehat{\Phi}^{i,j}_{p}\left[({\cal A}^{i,j}_{k,:})^{\top}\right]=\sum_{q=1}^{N}\widehat{\phi}^{i,j}_{pq}({\cal A}^{i,j}_{k,q})},$ via the Fourier bases such that
$\textstyle{\widehat{\phi}^{i,j}_{pq}({\cal A}^{i,j}_{k,q})=\sum_{m=1}^{G}a_{pqm}\cos{(m{\cal A}^{i,j}_{k,q})}+b_{pqm}\sin{(m{\cal A}^{i,j}_{k,q})},}$ (3)
where $m$ refers to the ${m^{\rm th}}$ grid point on a uniform grid of size $G$. The weights, [{${[\{[a_{pqm}],[b_{pqm}]\}_{m=1}^{G},W^{i,j}]}$ are updated in the backpropagation. If we use a Fourier basis in ${\widehat{\Phi}\in\mathbb{R}^{r\times N}}$, the parameter complexity of MHSA becomes ${O(2NrhGL)}.$ Regardless of the basis used, the linear projection takes $O(N^{2}rhL)$ FLOPs. For any Transformer, $L$ and $h$ are fixed. So, one may omit them from the complexity results. See FLOPs and memory requirement in Table 12.
Table 1: Performance of the best-performing Fourier KArAt models compared to the conventional vanilla ViT baselines. The best and the second-best Top-1 accuracies are given in red and blue, respectively. The $\downarrow$ and $\uparrow$ arrows indicate the relative loss and gain, respectively, compared to the base models.
@@
Table 2: Performance of Fourier KArAt fine-tuned on small datasets from ImageNet-1K pre-trained weights.
Model Acc.@1
SVHN Flowers 102 STL-10
ViT-Base 97.74 92.24 97.26
+ $G_{1}B$ 96.83 89.66 95.30
+ $G_{1}U$ 97.21 89.43 95.78
ViT-Small 97.48 91.46 96.09
+ $G_{3}B$ 97.04 89.67 95.26
+ $G_{3}U$ 97.11 90.08 95.45
ViT-Tiny 96.69 84.21 93.20
+ $G_{3}B$ 96.37 83.67 93.09
+ $G_{3}U$ 96.39 83.70 92.93
Blockwise and Universal operator configuration. We consider two configurations for updating the operator $\widehat{\Phi}^{i,j}$. (a) Blockwise: In this configuration, each encoder block learns the attention through $h$ distinct operators $\widehat{\Phi}$ for each of the $h$ heads, totaling $hL$ operators; see Figure 2 (a). Like the MHSA in vanilla ViTs, the blockwise configuration is designed to learn as many different data representations as possible. (b) Universal: This is motivated from the KAT [80]. In KAT, the MLP head is replaced with different variations of a KAN head—KAN and Group-KAN. In Group-KAN, the KAN layer shares rational base functions and their coefficients among the edges. Inspired by this, in our update configuration, all $L$ encoder blocks share the same $h$ operators; $\widehat{\Phi}^{i,j}=\widehat{\Phi}^{i}$ for $j=1,2,\ldots,L$; see Figure 2-(b). Rather than learning attention through $hL$ operators, this configuration only uses $h$ operators. Here, we share all learnable units and their parameters in each head across $L$ blocks. We postulate that the blockwise mode with more operators captures more nuances from the data. In contrast, the universal mode is suitable for learning simpler decision boundaries from the data. We also note that these two modes can be used with (2) as shown in Figure 1(b). Finally, Algorithm 2 in the Appendix gives pseudocode of applying Fourier-KArAt in the ViT encoder block. We provide more nuanced discussion about these two configurations in §B.3.6.
5 Empirical Study